# HMSViT: A Hierarchical Masked Self-Supervised Vision Transformer for Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis

### v2

@@ 3.2 Block Masked self-supervised learning
where the loss is computed only over the positions where $M$ indicates a mask (i.e., $1-M$ selects the masked regions).
@@ 3.3 Multi-scale decoder
where $\sigma(\cdot)$ denotes a convolution with activation functions to refine the aggregated representation.
@@ 4.1 Experimental Design
The experiments in this study are designed to evaluate the effectiveness and efficiency of our proposed self-supervised HMSViT across multiple tasks, focusing on diabetic peripheral neuropathy (DPN) diagnosis and segmentation of corneal nerve fibres. Three main experiments are conducted:
@@ 4.1.1 Task 1: Model Performance Evaluation
We evaluate HMSViT on two tasks—DPN diagnosis and corneal nerve fibre segmentation—by comparing three configurations (Tiny, Small, Base) with state-of-the-art models (Swin Transformer [11] and HiViT [12]).
@@
[…]
@@ 4.1.3 Task 3: Comparison with State-of-the-Art Methods for Segmentation
To comprehensively evaluate HMSViT’s segmentation performance, we compare it against several state-of-the-art transformer-based models, including standard Vision Transformers (ViTs), hierarchical ViTs, and self-supervised learning (SSL) approaches ((Table 3)). This evaluation spans diverse architectural paradigms, ranging from vanilla transformers to hierarchical and convolutional variants, ensuring a thorough comparison of model effectiveness.
@@
Table 3: Comparison of Vision Transformer architectures and Self-Supervised Learning approaches.
For a comprehensive performance assessment, we use evaluation metrics including mean Intersection over Union (mIoU), a widely adopted metric in segmentation tasks. In this experiment, we focused exclusively on corneal nerve segmentation results because precise nerve delineation is fundamental to both clinical assessment and the overall performance of our system. Accurate segmentation directly impacts the quantification of clinically relevant parameters—such as nerve fibre length and branch density—which are critical for diagnosing diabetic peripheral neuropathy. Moreover, most of the hierarchical models we compare are specifically designed for segmentation tasks, ensuring that our evaluation is both fair and directly aligned with the core functionality of these architectures.
@@ 4.2 Datasets
We employ two main datasets to train and evaluate HMSViT. This section details the acquisition, characteristics, and role of these datasets in our experiments. First, a larger, diverse collection of CCM images supports our self-supervised pretraining, enabling robust feature learning from unlabelled data. Second, a clinical dataset of corneal confocal microscopy (CCM) images—captured from both healthy volunteers and diabetic patients with or without peripheral neuropathy—is used for nerve segmentation and DPN diagnosis.
4.2.1 Datasets for self-supervised pretraining
In this work, a total of 6,426 eye CCM images from different sources were collected for block masked self-supervised learning pretrain (Table 4).
Dataset Description Count CORN-1 Nerve segmentation subset 1698 CORN-2 Image enhancement subset 688 CORN-3 Nerve tortuosity grading 403 CORN1500 Extended tortuosity grading 1500 UoM CCM Diabetes classification dataset 2137
Table 4: Datasets for Self-Supervised Learning.
4.2.2 Datasets for Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis
The dataset used in this study, provided by the Early Neuropathy Assessment (ENA) group from the University of Manchester, UK, consisted of images of the corneal sub basal nerve plexus. These images were sourced from healthy volunteers (HV) and participants with either prediabetes or diabetes, encompassing a total of 184 individuals.
Each participant had up to seven CCM images. The corneal confocal microscopy (CCM) images were obtained using an internationally recognized protocol developed by the ENA group. The imaging was conducted with the Rostock Corneal Module of a Heidelberg Retina Tomograph III confocal laser microscope (HRTII32-RCM, Heidelberg Engineering, Germany) at a resolution of 400×400 $\mu$m (384 × 384 pixels). The images were exported in BMP file format to ensure compatibility with the image analysis software.
@@ 4.2.2 Datasets for Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis
Figure 4: Confocal Corneal Microscopy (CCM) image showcasing expert-annotated nerve fibre segmentation and corneal nerve branch.
@@ 5.1 Model Performance Evaluation
Model Size Channels Blocks Param Nerve Segmentation Diagnostic classification mIoU Accuracy Accuracy Swin Tiny [96‑192‑384‑768] 12 29M 49.50% 96.00% 62.34% HiViT [96‑192‑384] 20M 48.00% 95.00% 61.25% HMSViT [96‑192‑384‑768] 28M 52.17% 98.50% 64.32% Swin Small [96‑192‑384‑768] 16 50M 51.00% 97.00% 64.32% HiViT [96‑192‑384] 38M 50.00% 96.50% 65.28% HMSViT [96‑192‑384‑768] 35M 53.23% 98.50% 67.60% Swin Base [128‑256‑512‑1024] 24 88M 53.00% 97.50% 68.42% HiViT [128‑256‑512] 67M 52.00% 97.00% 67.55% HMSViT [128‑256‑384‑512] 52M 55.76% 98.39% 70.40%
Table 5: Performance comparison of hierarchical transformer-based models for automated corneal nerve segmentation and diabetic neuropathy diagnosis.
@@
Table 6: Performance evaluation of various model configurations of HMSViT for corneal nerve fibre analysis on clinically relevant parameters, including corneal nerve fibre length (CNFL) and corneal nerve branch density (CNBD).
Diagnostic classification accuracy No Hierarchical With Hierarchical Supervised 65.60% 66.60% Self-supervised 69.40% 70.40% Corneal nerve segmentation performance Supervised 54.31% 57.28% Self-supervised 55.14% 61.34%
Table 7: Diagnostic classification accuracy and Corneal nerve segmentation performance for different training configurations, assessing the impact of hierarchical design and self-supervised learning.
We evaluated HMSViT on corneal nerve segmentation and diabetic neuropathy (DPN) diagnosis using three configurations (Tiny, Small, Base) and compared its performance with Swin Transformer [11] and HiViT [12].
Overall Performance: Table 5 shows that HMSViT consistently outperforms Swin and HiViT in both segmentation (mIoU) and diagnostic accuracy. For instance, in the Base configuration, HMSViT achieves a 55.76% mIoU versus 53.00% for Swin and 52.00% for HiViT. This superior performance can be attributed to its hierarchical multi-scale design and pooling-based token mixing, which effectively capture fine-grained details while preserving global context with fewer parameters.
[…]
Corneal Nerve fibre Analysis: Table 7 provides clinically relevant metrics for corneal nerve fibre length (CNFL) and branch density (CNBD). The Base configuration yields the lowest RMSE and SD values, indicating its superior ability to capture essential features.
Overall, HMSViT exhibits robust performance in both segmentation and diagnosis, leveraging efficient multi-scale feature extraction and advanced self-supervised learning for effective DPN screening.
Table 7 summarized the results of these settings in terms of classification accuracy and segmentation performance.
@@ 5.3 Comparison with State-of-the-Art Methods for Segmentation
Models Backbone Self-supervised Param mIoU (%) DPT[24] ViT None 97M 54.31 PoolFomer[40] HiViT None 77M 55.32 SegFormer[41] HiViT None 85M 57.39 MOCOV3+ViT ViT MOCOV3 86M 56.62 MAE+ViT ViT MAE 86M 57.08 MAE+HiViT HiViT MAE modified 73M 61.06 MixMIM[38] HiViT MAE modified 88M 60.86 Fast-iTPN[39] HiViT MAE modified 79M 61.25 Proposed method HMSViT MAE modified 68M 61.34
Table 8: Performance Comparison of Backbone Architectures and Self-Supervised Learning Methods in Nerve Segmentation.
In Table 8, we compared the performance and computational efficiency of state-of-the-art vision transformer-based segmentation models. The DPT model, built on a standard Vision Transformer architecture with 97M parameters, achieved an mIoU of 54.31%. In contrast, hierarchical models like PoolFormer (77M parameters) and SegFormer (85M parameters) showed modest improvements, reaching mIoU scores of 55.32% and 57.39%, respectively.
Notably, the application of self-supervised learning further enhanced performance. The models utilizing MAE-based techniques, such as MAE+ViT (86M parameters, mIoU: 57.08%) and MAE+HiViT (73M parameters, mIoU: 61.06%), benefited from the integration of a hierarchical design. Similarly, Fast-iTPN, incorporating a modified MAE framework with 79M parameters, achieves an mIoU of 61.25%.
Our proposed HMSViT achieved the highest mIoU of 61.34% with only 68M parameters, ensuring a strong balance between accuracy and efficiency. Its novel hierarchical multi-scale architecture and optimised self-supervised pretraining enhanced both fine-grained details and global context. These features made it well-suited for complex medical imaging tasks.

### v3

@@ 3.3 Block-masked self-supervised learning
where the loss is computed only over the positions where $M$ indicates a mask (i.e., 1M1-M selects the masked regions).
@@ 3.4 Multi-scale decoder
where $\sigma(\cdot)$ denotes a convolution with activation functions to refine the aggregated representation. The model is fine-tuned for this task using a combination of Dice loss and Binary Cross-Entropy (BCE) loss to handle class imbalance and produce accurate segmentation maps.
@@
This task is trained using a standard cross-entropy loss function to optimise classification accuracy. These task-specific loss functions are applied only during the fine-tuning stage, separately from the reconstruction loss used during self-supervised pretraining.
Figure 4: Architecture of the multi-scale decoder. For the segmentation task, hierarchical features from all stages are upsampled and fused to generate a high-resolution segmentation map. For the classification task, the global feature from the final stage is processed by a Multi-Layer Perceptron (MLP) to produce the diagnostic prediction.
@@ 4.1 Experimental Design
The experiments in this study are designed to comprehensively evaluate the effectiveness and efficiency of our proposed self-supervised HMSViT. Three main experiments are conducted: 1) Model Performance Evaluation, where we assess the model on DPN diagnosis and corneal nerve fibre segmentation; 2) Ablation Study, examining the impact of the hierarchical design and self-supervised learning; and 3) Comparison with State-of-the-Art Methods, benchmarking our segmentation results against existing techniques. Each of these tasks is discussed in detail below.
@@ 4.1.1 Task 1: Model Performance Evaluation
We evaluate HMSViT on two tasks—DPN diagnosis and corneal nerve fibre segmentation—by comparing three configurations (Tiny, Small, Base) with state-of-the-art hierarchical ViT based models (Swin Transformer [11] and HiViT [12]).
@@
    DPN Diagnosis Task: This task assesses the models’ ability to classify participants into three categories: Healthy Volunteers (HV), Diabetic Patients without Peripheral Neuropathy ($\text{PN}$-), and Diabetic Patients with Peripheral Neuropathy ($\text{PN}+$). To ensure a comprehensive evaluation, performance is measured at two levels:
    • [leftmargin=*, topsep=0pt, partopsep=0pt, itemsep=0pt]
    • Patient-Level Diagnosis: To mitigate intra-subject variability and improve diagnostic robustness, predictions from the 4–8 images per participant are aggregated via majority voting [40]. Majority voting reflects clinical practice, where diagnosis is made per patient rather than per image. The resulting patient-level accuracy serves as the primary metric for comparing HMSViT against other state-of-the-art hierarchical models.
    • Image-Level Diagnosis: To evaluate the intrinsic classification capability of our model, we also report image-level metrics—including precision, recall, F1-score, and accuracy—for each of the three HMSViT configurations (Tiny, Small, and Base).
@@ 4.1.3 Task 3: Comparison with State-of-the-Art Methods for Segmentation
To comprehensively evaluate HMSViT’s segmentation performance, we benchmark it against several state-of-the-art Vision Transformer-based models, including baseline ViTs, hierarchical ViTs, and self-supervised learning (SSL) approaches. These existing methods are summarised in Table 3. Our evaluation spans diverse architectural paradigms, ranging from vanilla transformers to hierarchical and convolutional variants, ensuring a thorough comparison of model effectiveness.
@@
Table 3: Summary of representative Vision Transformer-based models and self-supervised learning approaches used for medical image segmentation. The models are classified into baseline ViTs, hierarchical architectures, and self-supervised frameworks, highlighting their core design strategies and contributions to efficient feature learning.
For a comprehensive performance assessment, we use evaluation metrics including mean Intersection over Union (mIoU), a widely adopted metric in segmentation tasks. In this experiment, we focus exclusively on corneal nerve segmentation results because precise nerve delineation is fundamental to both clinical assessment and the overall performance of our system. Accurate segmentation directly impacts the quantification of clinically relevant parameters—such as nerve fibre length and branch density—which are critical for diagnosing DPN. Moreover, most of the hierarchical models we compare are specifically designed for segmentation tasks, ensuring that our evaluation is both fair and directly aligned with the core functionality of these architectures.
@@ 4.2 Datasets
Our study leverages two distinct data sources to train and evaluate HMSViT: a public dataset and a private clinical cohort. The first is the public CORN CCM dataset [44], an aggregation of four subsets designed for various tasks including nerve segmentation, image enhancement, and tortuosity grading. This dataset provides a diverse collection of CCM images that enriches our self-supervised pretraining phase. The second is a private clinical UoM CCM dataset from the Early Neuropathy Assessment (ENA) group at the University of Manchester, UK, which includes CCM images from 318 participants.
These two sources serve distinct but complementary roles: self-supervised pretraining and two downstream tasks including corneal nerve segmentation and DPN diagnosis. To facilitate robust self-supervised learning (SSL), we aggregated a large pretraining dataset of 6,426 corneal confocal microscopy (CCM) images from multiple sources. As detailed in Table 4, this collection includes the four subsets of the public CORN dataset [44] and the clinical CCM Dataset (UOM CCM) mentioned above. For pretraining, all existing labels were disregarded, and only the raw image data was used to enable the model to learn fundamental feature representations.
Dataset Description Count
CORN-1 Nerve segmentation subset 1698
CORN-2 Image enhancement subset 688
CORN-3 Nerve tortuosity grading 403
CORN1500 Extended tortuosity grading 1500
UoM CCM Diabetes classification dataset 2137
Table 4: Datasets aggregated for self-supervised pretraining.
@@
The participants were categorised into three groups: Healthy Volunteers (HV): 112 patients (35.2%), no diabetes or neuropathy. Diabetic Patients without Neuropathy ($\text{PN}$-): 98 patients (30.8%), diabetes duration $\geq$ 5 years. Diabetic Patients with Neuropathy ($\text{PN}+$): 108 patients (34.0%), confirmed DPN symptoms. The demographic and clinical characteristics of the study participants are summarised in Table 5.
Table 5: Demographic and clinical characteristics of the study participants. Values are presented as Mean $\pm$ Standard Deviation (SD) or Count.
Variable Control with DPN (PN+) without DPN (PN-)
Duration of diabetes (years) 0 $\pm$ 0 45.54 $\pm$ 11.96 23.5 $\pm$ 13.74
Age 39.64 $\pm$ 14.15 60.79 $\pm$ 11.40 41.85 $\pm$ 14.46
Gender (Female/Male) 56.4%/43.6% 52.8%/47.2% 44.6%/55.4%
BMI ($kg/m^{2}$) 25.64 $\pm$ 4.74 27.55 $\pm$ 3.90 26.22 $\pm$ 4.56
HbA1c 5.56 $\pm$ 0.34 8.27 $\pm$ 1.34 8.34 $\pm$ 1.45
IFCC (mmol/mol) 36.34 $\pm$ 5.51 66.60 $\pm$ 14.25 67.60 $\pm$ 15.79
The manual annotations of the nerve fibre centerlines were traced by an ophthalmologist using the open-source software ImageJ [45]. Figure 5 illustrates the corneal nerve fibre analysis, displaying the original CCM images, manual annotations, and automated nerve fibre segmentation analysis. In this analysis, longer nerve fibres with larger diameters are identified as main trunks, while fibres originating from these trunks are classified as nerve branches. This classification, performed using ACCMetrics [46], forms the basis for calculating CNFL and CNBD.
Figure 5: Visualisation of corneal nerve fibre analysis from corneal confocal microscopy (CCM) images. Top row: original corneal nerve images from three subjects. Middle row: ground-truth manual annotations (cyan traces). Bottom row: automated nerve fibre segmentation showing main nerves (red) and branches (blue), with detected branching points (green dots).
4.3 Implementation Details
All experiments were conducted on a workstation equipped with an NVIDIA RTX A6000 GPU and an Intel Xeon W-2255 CPU, using PyTorch 2.1 with CUDA 12 and mixed-precision training.
Self-supervised pretraining: We employed a block-masked reconstruction strategy with a patch size of $4\times 4$, and blocks composed of $4\times 4$ patches. Sensitivity analysis (ratios 50%–90%) identified 75% as the optimal masking ratio (yielding 85.6% diagnostic accuracy and 61.34% mIoU), balancing reconstruction difficulty with semantic retention—lower ratios made tasks too trivial, while higher ratios removed critical context. Models were trained for 400 epochs using the AdamW optimiser (learning rate 1.5e-4) with a cosine decay schedule following a 10-epoch warm-up.
Downstream tasks training: Models were fine-tuned for 160 epochs using a combined loss function and the same cosine decay schedule.
To ensure robust and unbiased evaluation, we implemented a patient-level 5-fold cross-validation scheme. The 318 participants were randomly partitioned into five folds, with all images from a single participant confined to the same fold to prevent data leakage. In each cross-validation run, three folds were used for training, one for validation, and one for testing. The final reported metrics are the average performance across the five test folds. The source code for this project will be made publicly available at https://gitlab.com/han-research/hmsvit.
@@ 5.1 Model Performance Evaluation
Model Size Channels Blocks Param Nerve Segmentation Diagnostic Accuracy mIoU(%) Accuracy Image-Level Patient-Level Swin Tiny [96‑192‑384‑768] 12 29M 54.45 ± 0.72 0.960 ± 0.005 0.623 ± 0.015 0.762 ± 0.013 HiViT Tiny [96‑192‑384] 20M 52.80 ± 0.78 0.950 ± 0.006 0.613 ± 0.016 0.753 ± 0.014 HMSViT Tiny [96‑192‑384‑768] 28M 57.39 ± 0.48 0.985 ± 0.003 0.643 ± 0.010 0.785 ± 0.009 Swin Small [96‑192‑384‑768] 16 50M 56.10 ± 0.65 0.970 ± 0.005 0.643 ± 0.013 0.794 ± 0.012 HiViT Small [96‑192‑384] 38M 55.00 ± 0.70 0.965 ± 0.006 0.653 ± 0.014 0.802 ± 0.013 HMSViT Small [96‑192‑384‑768] 35M 58.55 ± 0.45 0.985 ± 0.003 0.676 ± 0.009 0.832 ± 0.008 Swin Base [128‑256‑512‑1024] 24 88M 58.30 ± 0.68 0.975 ± 0.005 0.684 ± 0.012 0.823 ± 0.010 HiViT Base [128‑256‑512] 67M 57.20 ± 0.73 0.970 ± 0.006 0.676 ± 0.013 0.817 ± 0.012 HMSViT Base [128‑256‑384‑512] 52M 61.34 ± 0.40 0.984 ± 0.003 0.704 ± 0.008 0.856 ± 0.007
Table 6: Performance comparison of the proposed HMSViT with hierarchical vision transformers Swin [11] and HiViT [12]. For the Tiny, Small, and Base variants, we report each model’s backbone configuration (channels, blocks) and parameter count (Params), together with performance on corneal nerve segmentation and diabetic neuropathy diagnosis. Values are mean ± SD over 5-fold cross-validation.
Figure 6: Performance gains of HMSViT over Swin and HiViT across Tiny, Small, and Base configurations in patient-level diagnostic accuracy and segmentation mIoU.
Class Tiny Small Base Precision Recall F1-score Precision Recall F1-score Precision Recall F1-score HV 0.730 0.790 0.760 0.695 0.802 0.745 0.812 0.758 0.784 PN+ 0.570 0.530 0.550 0.667 0.667 0.667 0.615 0.667 0.640 PN- 0.550 0.510 0.530 0.650 0.527 0.582 0.645 0.662 0.653 Accuracy 0.643 0.676 0.704
Table 7: Detailed image-level diagnostic classification performance of HMSViT across Tiny, Small, and Base configurations. The table presents precision, recall, and F1-score for each class: Healthy Volunteers (HV), Diabetic Patients with Neuropathy (PN+), and Diabetic Patients without Neuropathy (PN-).
We evaluated HMSViT on corneal nerve segmentation and diabetic neuropathy (DPN) diagnosis using three configurations (Tiny, Small, Base) and compared its performance with state-of-the-art hierarchical models, Swin Transformer [11] and HiViT [12].
Overall Performance: As detailed in Table 6 and illustrated in Figure 6, HMSViT consistently outperforms both Swin Transformer and HiViT across all configurations in nerve segmentation and diagnostic accuracy. Notably, the HMSViT-Base model achieves a patient-level diagnostic accuracy of 85.6% and a segmentation mIoU of 61.34%, surpassing Swin-Base (82.3% accuracy, 58.30% mIoU) and HiViT-Base (81.7% accuracy, 57.20% mIoU). This performance advantage is achieved with substantially fewer parameters (52M) than Swin (88M) and HiViT (67M), underscoring the efficiency of the proposed pooling-based hierarchical architecture and the block-masked self-supervised pretraining strategy.
Diagnostic Classification Performance: Table 7 provides a detailed breakdown of the image-level classification performance. HMSViT-Base demonstrates strong and balanced metrics across all classes, achieving an F1-score of 0.784 for Healthy Volunteers (HV), 0.640 for PN+, and 0.653 for PN-. The overall image-level accuracy for HMSViT-Base is 70.40%, which is higher than Swin-Base (68.42%) and HiViT-Base (67.55%) as shown in Table 6. This indicates that the features learned through our self-supervised approach are highly discriminative for identifying subtle neuropathic changes.
Corneal Nerve Fibre Analysis: In addition to segmentation and classification, we assessed the model’s ability to quantify clinically relevant nerve morphology parameters. Table 9 shows the Root Mean Square Error (RMSE) and Standard Deviation (SD) for Corneal Nerve Fibre Length (CNFL) and Corneal Nerve Branch Density (CNBD). The HMSViT-Base configuration yields the lowest RMSE (326.002 for CNFL, 4.1603 for CNBD) and SD values, indicating its superior precision in quantifying nerve structures compared to the Tiny and Small variants. This highlights the model’s potential for reliable clinical biomarker extraction.
In summary, HMSViT demonstrates robust and superior performance in both segmentation and diagnosis, leveraging efficient multi-scale feature extraction and advanced self-supervised learning to deliver a powerful tool for DPN screening.
@@ 5.2 Ablation Study to Evaluate the Impact of Hierarchical Design and Self-Supervised Learning
Table 8: Performance evaluation of hierarchical multi-scale Vision Transformer (HMSViT) variants (Tiny, Small, Base) on corneal nerve biomarker prediction. The table reports Root Mean Square Error (RMSE) and standard deviation (SD) for Corneal Nerve Fibre Length (CNFL) and Corneal Nerve Branch Density (CNBD). The Base model achieves the lowest error and variability for both biomarkers, indicating improved predictive performance with increased model capacity.
Diagnostic classification accuracy No Hierarchical With Hierarchical Supervised 0.656 0.666 Self-supervised 0.694 0.704 Corneal nerve segmentation performance (%) Supervised 54.31 57.28 Self-supervised 55.14 61.34
Table 9: Performance comparison of supervised and self-supervised learning approaches for diagnostic classification and corneal nerve segmentation, with and without hierarchical modelling. Self-supervised methods outperform supervised ones across both tasks, with notable improvements when hierarchical modelling is incorporated, particularly in corneal nerve segmentation accuracy.
Table 9 summarised the results of these settings in terms of classification accuracy and segmentation performance.
@@ 5.3 Comparison with State-of-the-Art Methods for Segmentation
Models Backbone Self-supervised Param Inf. Time (ms) mIoU (%) DPT[24] ViT None 97M 22.8 54.31 PoolFormer[41] HiViT None 77M 17.1 55.32 SegFormer[42] HiViT None 85M 19.5 57.39 MOCOV3[43]+ViT ViT MOCOV3 86M 20.1 56.62 MAE[35] +ViT ViT MAE 86M 19.8 57.08 MAE+HiViT[12] HiViT MAE modified 73M 16.5 61.06 MixMIM[38] HiViT MAE modified 88M 20.7 60.86 Fast-iTPN[39] HiViT MAE modified 79M 17.9 61.25 Proposed method HMSViT MAE modified 68M 15.2 61.34
Table 10: Benchmarking HMSViT against state-of-the-art segmentation models. Our proposed method achieves the highest mIoU with the fewest parameters and fastest inference time, demonstrating a superior balance of accuracy and efficiency.
In Table 10, we compared the performance and computational efficiency of state-of-the-art vision transformer-based segmentation models. The DPT model, built on a standard Vision Transformer architecture with 97M parameters, achieved an mIoU of 54.31% with an inference time of 22.8 ms. In contrast, hierarchical models like PoolFormer (77M parameters, 17.1 ms) and SegFormer (85M parameters, 19.5 ms) showed modest improvements, reaching mIoU scores of 55.32% and 57.39%, respectively.
Notably, the application of self-supervised learning further enhanced performance. The models utilising MAE-based techniques, such as MAE+ViT (86M parameters, 19.8 ms, mIoU: 57.08%) and MAE+HiViT (73M parameters, 16.5 ms, mIoU: 61.06%), benefited from the integration of a hierarchical design. Similarly, Fast-iTPN, incorporating a modified MAE framework with 79M parameters, achieves an mIoU of 61.25% in 17.9 ms.